OpenStudy (anonymous):

Complete the square and find the x and y intercepts. y=-2x^2+3x+4

terenzreignz (terenzreignz):

To complete the square, you first make the x^2 "alone", that is to say, it must have 1 as its coefficient :)

OpenStudy (anonymous):

I got -2(x-3/4)+41/9

OpenStudy (anonymous):

y-intercept = 4 x-intercept=2.35 and -0.851

terenzreignz (terenzreignz):

Nicely done, more quickly than I did, :D

OpenStudy (anonymous):

@sauravshakya could you show me your work?

OpenStudy (anonymous):

for y-intercept, x=0 so, y =-2*(0)^2+3*(0)+4=4 thus,4 is the y-intercept for x intercept, y=0 so, 0=-2x^2+3x+4 or,2x^2-3x-4=0 Now use any method to solve the above eqn to get the value of x. And this values of x will be x-intercept.

OpenStudy (phi):

complete the square \[ -2(x^2 + -\frac{3}{2}x -2) = y \] multiply the coeff of x by 1/2 and then square it: -3/2 * 1/2 = -3/4 squared: 9/16 add and subtract 9/16 to the stuff inside the parens (9/16-9/16=0 so we are not changing its value): \[ -2(x^2 + -\frac{3}{2}x +\frac{9}{16}-\frac{9}{16}-2) = y \] group the first 3 terms inside the parens: \[ -2( (x^2 + -\frac{3}{2}x +\frac{9}{16})-\frac{9}{16}-2) = y \] the grouped terms are a perfect square: \[ -2( (x - \frac{3}{4})^2-\frac{9}{16}-2) = y \] combine the last 2 terms in side the parens: \[ -2( (x - \frac{3}{4})^2-\frac{41}{16}) = y \]

OpenStudy (phi):

to find the x intercepts, set y equal to zero \[ -2( (x - \frac{3}{4})^2-\frac{41}{16}) = 0 \] divide through by -2 to simplify \[ (x - \frac{3}{4})^2-\frac{41}{16} = 0 \] move the constant to the right side \[ (x - \frac{3}{4})^2=\frac{41}{16} \] take the square root of both sides \[ x - \frac{3}{4}=\frac{\pm\sqrt{41}}{4} \] add 3/4 to both sides \[x = \frac{1}{4}(3\pm\sqrt{41}) \]